# tidychangepoint: a unified framework for analyzing changepoint detection in univariate time series

### v2

@@ 1 Introduction
1.1 Motivating example
While global temperatures continue to rise, climate scientists continue to face baffling skepticism from some. How do we know that the increase we observe are not just anecdotal? How do we know that the recent pattern of rising temperatures is not simply the latest round of normal fluctuation?
@@ 1.1 Motivating example
For example, the CET data set collected by Parker, Legg, and Folland (1992) and repackaged in tidychangepoint (Baumer et al. 2024) lists the mean annual temperature in degrees Celsius from 1659 to 2024, as measured in Central England and shown in Figure 1. When did the temperature start to rise? How can we distinguish natural variation in temperature (around a static mean) from a meaningful change in that mean (caused by humans)?
@@
Changepoint detection analysis is the study of such changepoints and the algorithms and models that determine them. changepoint (Killick 2024) is a popular R package that implements the PELT algorithm (Killick, Fearnhead, and Eckley 2012) for changepoint detection. We can apply this algorithm to our temperature data using the cpt.meanvar() function.
@@
In this case, PELT finds 6 changepoints, listed below and illustrated in Figure 2.
index(CET)[cet_cpt@cpts] |>
  format("%Y")
[1] "1713" "1715" "1925" "2002" "2005" "2024"
While PELT has many virtues, there are many alternative changepoint detection algorithms. Another algorithm is Wild Binary Segmentation, provided by the wbs package (Baranowski and Fryzlewicz 2024).
@@
index(CET)[sort(cet_wbs\$cpt\$cpt.ic\$mbic.penalty)] |>
  format("%Y")
[1] "1701" "1988"
This algorithms finds 2 changepoints, as seen in Figure 2.
(a) PELT. Plot created by plot.cpt().
(b) WBS. Plot created by plot.wbs().
Figure 2: Changepoints found in temperatue data by two algorithms.
Another promising approach to determining changepoints is to use genetic algorithms that iterate over many possible changepoint sets. Shi et al. (2022) explored the use of a variety of different parametric models using genetic algorithms and a previous version of the CET data set. This analysis leveraged the GA package, which is not specific to changepoint detection, and thus each model that Shi et al. (2022) wanted to test required writing a different set of code, the complexity of which prevents us from reproducing it here. Using a trendshift model (see Section 5), they found the optimal changepoints to be 1700, 1739, and 1988.
There is no optimal solution to the Changepoint Detection Problem (CDP, defined formally in Section 1.2), and in this case, the three algorithms returned different changepoint sets, even when applied to the same data set. Furthermore—and more germaine for us—the objects returned by changepoint detection functions from changepoint, wbs, and GA are all of different types, have difffent structures, contain different information, and can be used with a variety of different functions. This heterogeneity in user interface and package architecture is widespread across roughly a dozen other changepoint detection packages we found. No standardization exists.
While we do not claim to offer solutions to the CDP itself, tidychangepoint solves the software heterogeneity problem by providing a unified interface with built-in support for genetic algorithms for changepoint detection. Comparing algorithms and/or models for changepoint detection in tidychangepoint can be as easy as iterating over a list of specifications.
@@
  shi = segment(CET, method = "ga-shi"maxiter = 1000run = 100)
Once the algorithms have been run, we can easily extract the sets of changepoints for comparison using the changepoints() function or compare the results visually using plot() (see Figure 3). Each panel of Figure 3 shows the original time series (black line) segmented by the changepoints determined by the PELT algorithm (vertical dotted lines). The horizontal red lines indicate the mean and 1.96 times the standard deviation within each region, as per the summary produced by tidy() (see Section 3.2).
@@
  lapply(plot, use_time_index = TRUEylab = "Degrees Celsius"|>
  patchwork::wrap_plots()
Figure 3: Comparison of four different attempts to detect changepoints in the Central England temperature data. Plots created by plot.tidycpt().
cet_objs |>
  lapply(changepoints, use_labels = TRUE|>
@@ 1.2 Mathematical problem formulation
Let $y=\{y_{1},\ldots,y_{n}\}$ be a series of observations made over time points $t=1,\ldots,n$. Define $\tau=\{\tau_{1},\ldots,\tau_{m}\}\subseteq y$ to be a subset of the original observations known as changepoints, for some $0\leq m\leq n$. The $m$ changepoints divide the time series into $m+1$ regions, with the idea that the behavior of the time series within each region does not change, and the changepoints represent points in time at which the behavior of the underlying system changes in some meaningful way. If we adopt the convention that $\tau_{0}=1$ and $\tau_{m+1}=n+1$ then the union of the set of half-open intervals $[\tau_{j},\tau_{j+1})$ for $0\leq j\leq m$ contains the time series $y$.
@@
As there as $2^{n}$ possible changepoint sets, the Changepoint Detection Problem (CDP)—which is the problem of finding the optimal changepoint set (according to some metric)—has attracted considerable attention in the literature. Aminikhanghahi and Cook (2017) provide an excellent survey of relevant results. In unsupervised settings, there is no known true changepoint set, and various models, metrics, and algorithms have been proposed, some of which are in widespread use. We briefly highlight the previous work most relevant to ours in Section 1.3, and review existing software implementations of these algorithms in Section 1.4. In any case, all attempts to solve the CDP considered in this paper involve three essential components: 1) a parametric model $M(y|\theta)$ that describes the data-generating process of the time series; 2) an algorithm $A$ to search the exponentially large space of possible changepoint sets; and 3) a metric $f:(\tau,M)\rightarrow\mathbb{R}$ (a penalized objective function) for evaluating the suitability of any given changepoint set. We define $\tau_{f,M,\theta}^{*}$ to be the changepoint set that minimizes the value of the penalized objective function $f$ on the model $M$ with parameters $\theta$ over the space of all possible changepoint sets. That is,
$\tau_{f,M,\theta_{\tau}}^{*}=\min_{\tau\subseteq y}f(\tau,M(y|\theta_{\tau}))$
@@
Note that the value of the parameters $\theta$ depend on the changepoint set $\tau$, so we may write $\theta_{\tau}$.
@@
For example, perhaps the simplest and most common parametric model for changepoint detection is the normal (Gaussian) meanshift model, in which the time series is assumed to be generated by random variables $Y_{j}\sim N(\mu_{j},\sigma^{2})$ for each region $11\leq j\leq m+$ 1. Thus the series mean $\mu_{j}$ is constant within each of the $m+1$ regions, but varies across the regions, while the variance $\sigma^{2}$ is constant over the entire time series. We fit the normal meanshift model using Maximium Likelihood Estimation (MLE), resulting in the estimated parameters $\hat{\theta}=\{\hat{\sigma}^{2},\hat{\mu}_{0},\ldots,\hat{\mu}_{m}\}$. If we choose the Bayesian Information Criterion (BIC) (Schwarz 1978) as our penalty, then the penalized objective function becomes:
@@
where $k_{M}(\tau)$ is the number of parameters estimated by the model (in this case, $m+2$) and $L_{M}(y|\hat{\theta}_{\tau})$ is the likelihood function for $M$, which in this case is (see Li and Lund (2012) for details):
@@
More generally, given a time series $y$, a parametric model $M(y|\theta)$, and a penalized objective function $f$, the goal of any changepoint detection algorithm $A$ is to search the exponentially large space of possible changepoints sets for the one $\tau^{*}$ that minimizes the value of $f(\tau,M(y|\hat{\theta}_{\tau}))$. Of course, in the process it must also estimate $\hat{\theta}$.
1.3 Existing changepoint detection algorithms
@@
The Pruned Exact Linear Time (PELT) algorithm developed by Killick, Fearnhead, and Eckley (2012) builds on a dynamic programming algorithm from Jackson et al. (2005) to offer a linear time, exact algorithm for changepoint detection. PELT offers order of magnitude improvements in computational cost (i.e., $O(n)$ vs. the $O(n\log{n})$ performance of the binary segmentation algorithm (Scott and Knott 1974) and the $O(n^{3})$ performance of the segmented neighborhood algorithm (Auger and Lawrence 1989)), and is capable of optimizing over different parametric models (e.g., meanvar, meanshift, etc.) and different penalized objective functions (e.g., BIC, mBIC, etc.). To guarantee an optimal solution, PELT requires that the value of the penalized objective function decreases (or increases by an amount bounded by a known constant) with each additional changepoint. Moreover, its running time can be quadratic (i.e., $O(n^{2})$) unless the number of changepoints grows linearly with the length of the time series. While these assumptions are mild, not all possible combinations of time series, models, and penalty functions will meet them.
@@
Li and Lund (2012) illustrate how a genetic algorithm can provide good results by employing Darwinian principles (e.g., survival of the fittest) to search the space of all possible changepoint sets intelligently. For a given parametric model and penalized objective function, a genetic algorithm starts with an initial “generation” of (usually randomly generated) candidate changepoint sets, evaluates them all using the penalized objective function, then probabilistically “mates” the most promising candidate changepoint sets to produce a new generation of the same size. This process is repeated until a stopping criteria is met, and the single best (fittest) changepoint set is returned. While genetic algorithms are neither exact, nor deterministic, nor computationally efficient, they are more easily generalizable than PELT, and can return competitive results in practice (Shi et al. 2022). Li and Lund (2012) also develop the notion of the Minimum Descriptive Length (MDL) as a penalty function based on information theory.
@@
Bai and Perron (1998) use linear regression models fit by least squares to detect changepoints. In particular, they construct a framework for testing the alternative hypothesis that there is one additional changepoint. Barry and Hartigan (1993) employ Bayesian techniques for changepoint detection. Cho and Fryzlewicz (2015) propose the Sparsified Binary Segmentation (SBS) algorithm, which uses CUSUM statistics to analyze high dimensional time series. This algorithm uses the concept of a threshold above which statistics are kept—an idea that resurfaces in Suárez-Sierra, Coen, and Taimal (2023). Cho (2016) explores the utility of a double CUSUM statistic for panel data. Hocking et al. (2013) use machine learning techniques to determine the optimal number of changepoints.
@@ 1.4 Existing changepoint detection R packages
Likely the most prominent R package for changepoint detection is changepoint (Killick 2024; Killick and Eckley 2014), which implements the algorithms PELT, Binary Segmentation, and Segmented Neighborhood. PELT employs one of three different models: meanshift, meanvar, and varshift (see Section 5). For the meanvar model (the most versatile) the data-generating model can follow a Normal, Poisson, Gamma, or Exponential distribution. Penalty functions (see Section 6) implemented by PELT include AIC, BIC, and MBIC, but not MDL. EnvCpt [R-EnvCpt] is another package that wraps changepoint functionality.
The wbs (Baranowski and Fryzlewicz 2024) package implements the Wild Binary Segmentation and standard Binary Segmentation algorithms, the first of which is incorporated into tidychangepoint. The ggchangepoint (Yu 2022) package provides some wrapper functions and graphical tools for changepoint detection analysis that are similar to ours, but more limited in scope and usability. Nonparametric methods for changepoint detection are present in the ecp (James, Zhang, and Matteson 2024) and changepoint.np (Haynes and Killick 2022) packages. We consider these beyond our scope at this time, which is focused on parameteric models for changepoint detection. A Bayesian approach is implemented in the bcp (X. Wang, Erdman, and Emerson 2018; Erdman and Emerson 2008) package, which is no longer available on CRAN. The qcc (Scrucca 2017) package is more broadly focused on quality control and statistical process control, neither of which are specific to changepoint detection. The mcp package focuses on regression-based, Bayesian methods. There are many other changepoint packages, all of which have their strengths and limitations. Please see the “Comparison to other packages” vignette in the mcp documentation for a thorough comparison.

### v3

# captionstableheading

@@ 1 Introduction
1.1 Motivation
While global temperatures continue to rise, climate scientists continue to face baffling skepticism from some. How do we know that the increases we observe are not just anecdotal? How do we know that the recent pattern of rising temperatures is not simply the latest round of normal fluctuation?
@@ 1.1 Motivation
Changepoint detection analysis is the study of such changepoints and the algorithms and models that determine them. Unfortunately, there is no known general solution to the Changepoint Detection Problem (CDP, defined formally in Section 1.4). Furthermore—and more germaine for us—the objects returned by changepoint detection functions from various R packages (see Section 1.6) are all of different types, have different structures, contain different information, and can be used with a variety of different functions. This heterogeneity in user interface and package architecture is widespread across roughly a dozen other changepoint detection packages we found (see Section 1.6). No standardization exists.
While we do not claim to offer solutions to the CDP itself, tidychangepoint solves the software heterogeneity problem by providing a unified interface with built-in support for genetic algorithms for changepoint detection. Comparing algorithms and/or models for changepoint detection in tidychangepoint can be as easy as iterating over a list of specifications (see Section 1.3).
1.2 An example illustrating the status quo
For example, the CET data set collected by Parker, Legg, and Folland (1992) and repackaged in tidychangepoint (Baumer et al. 2025) lists the mean annual temperature in degrees Celsius from 1659 to 2024, as measured in Central England and shown in Figure 1. When did the temperature start to rise? How can we distinguish natural variation in temperature (around a static mean) from a meaningful change in that mean (caused by humans)?
@@ 1.2 An example illustrating the status quo
Suppose that you are interested in determining changepoints for the CET data set, but you don’t know much about the inner workings of changepoint detection algorithms. You decide to simply fit a bunch of off-the-shelf changepoint detection models and compare the results. Some googling leads you to four packages: strucchange (Zeileis et al. 2024), segmented (V. M. R. Muggeo 2025), changepoint (Killick 2024), and wbs (Baranowski and Fryzlewicz 2024).
First, you fit four models. Note that because CET is an xts object, sometimes (but not always), you have to cast it to a ts object.
@@
cet_bkp <- strucchange::breakpoints(CET ~ 1)
cet_seg <- segmented::segmented(as.ts(CET))
Next, you seek to recover the changepoints as a numeric vector. These are the indices of the time series (in this case, each corresponds to a year since 1659). Note that each object (some of which are S3 and some of which are S4) stores the changepoints in different locations, some of which are not easy to find.
cet_cpt@cpts
[1]  55  57 267 344 347 366
sort(cet_wbs\$cpt\$cpt.ic\$mbic.penalty)
[1]  29  43  81  82 234 330
cet_bkp\$breakpoints
[1]  55 252 312
cet_seg\$psi[ , 2]
[1] 321.788
Finally, you plot each object using plot(). Figure 11 in our Appendix shows the resulting plots, some of which add changepoints and model estimates to the original time series, and others of which show diagnostic plots.
While you were able to conduct your analysis, you had to learn the ins-and-outs of four different packages, all of which were designed in their own way.
1.3 An example using tidychangepoint
tidychangepoint makes the comparisons in the previous section simpler by employing a unified syntax and output object. With tidychangepoint, it is possible to iterate analyses over lists of specifications.
We demonstrate this approach using the same four models as above, plus a null model that finds no changepoints, and a genetic algorithm that Shi et al. (2022) use to explore a variety of different parametric models using genetic algorithms and a previous version of the CET data set. [^1] Note that we have set the maximum number of iterations (generations) to just 1000. To replicate the exact results of Shi et al. (2022) we would need a much larger value. As noted in Section 7.1, the values in the underlying data set have also changed. However, we note that even with different data and many fewer generations, we identified all three of the changepoints found by Shi et al. (2022) to within a year (along with three others). Using a trendshift model (see Section 5), they identified the years 1700, 1739, and 1988 as changepoints. [^2] Because this analysis leveraged the GA (Scrucca 2024) package, which is not specific to changepoint detection, each model that Shi et al. (2022) wanted to test required writing a different set of code, the complexity of which prevents us from reproducing it here..
@@ 1.3 An example using tidychangepoint
bkp = segment(CET, method = "strucchange"),
seg = segment(CET, method = "segmented"),
shi = segment(CET, method = "ga-shi", maxiter = 1000, run = 100)
Once the algorithms have been run, we can easily extract the sets of changepoints for comparison using the changepoints() function and convert them to human-interpretable years with the as_year() function.
@@
lapply(changepoints, use_labels = TRUE) |>
@@
We compare the results visually using plot() (see Figure 2), which now always situates the model estimates in the domain of the original time series. Each panel of Figure 2 shows the original time series (black line) segmented by the changepoints determined by the various algorithms (vertical dotted lines). The horizontal red lines indicate the mean and 1.96 times the standard deviation within each region, as per the summary produced by tidy() (see Section 3.2).
cet_objs |>
lapply(plot, use_time_index = TRUE, ylab = "Degrees Celsius") |>
patchwork::wrap_plots()
Figure 2: Comparison of four different attempts to detect changepoints in the Central England temperature data. Note that the model in the lower-right panel includes an autoregressive parameter, and thus the model estimates vary across years within regions. All plots created by plot.tidycpt().
@@ 1.4 Mathematical problem formulation
Let $y=\{y_{1},\ldots,y_{n}\}\in\mathbb{R}^{n}$ be a discrete series of observations made over an integer-valued index of time points $t=1,\ldots,n$. Define $\tau=\{\tau_{1},\ldots,\tau_{m}\}\in[1,n]^{m}$ to be a set of time point indexes known as changepoints, for some $0\leq m\leq n$. The $m$ changepoints divide the time series into $m+1$ regions, with the idea that the behavior of the time series within each region does not change, and the changepoints represent points in time at which the behavior of the underlying system changes in some meaningful way. If we adopt the convention that $\tau_{0}=1$ and $\tau_{m+1}=n+1$ then the union of the set of half-open intervals $\tau_{j},\tau_{j+1})$ for $0\leq j\leq m$ covers the time domain of the series $y$.
@@
Even if the changepoints are assumed to come from the discrete set $t$, there are $2^{n}$ possible changepoint sets, and the Changepoint Detection Problem (CDP)—which is the problem of finding the optimal changepoint set (according to some metric)—has attracted considerable attention in the literature. Aminikhanghahi and Cook (2017) provide an excellent survey of relevant results. In the typical unsupervised setting in which the true changepoints are not known, various models, metrics, and algorithms to find the optimal changepoint set have been proposed, some of which are in widespread use. We briefly highlight the previous work most relevant to ours in Section 1.5, and review existing software implementations of these algorithms in Section 1.6. In any case, all attempts to solve the CDP considered in this paper involve three essential components: 1) a parametric model $M(y|\theta)$ that describes the data-generating process of the time series; 2) an algorithm $A$ to search the exponentially large space of possible changepoint sets; and 3) a metric $f:(\tau,M)\rightarrow\mathbb{R}$ (a penalized objective function) for evaluating the suitability of any given changepoint set. We define $\tau_{f,M,\theta}^{*}$ to be the changepoint set that minimizes the value of the penalized objective function $f$ on the model $M$ with parameters $\theta$ over the space of all possible changepoint sets. That is,
$\tau_{f,M,\theta_{\tau}}^{*}=\arg\min_{\tau\in[1,n]^{m}}f(\tau,M(y|\theta_{\tau}))$
@@
The value of the penalized objective function $f(\tau,M(y|\hat{\theta}_{\tau}))$ is often referred to as the fitness of the changepoint set $\tau$ and the model $M$ with estimated parameters $\hat{\theta}_{\tau}$ to the time series $y$. Note that the value of the parameters $\theta$ depend on the changepoint set $\tau$, so we may write $\theta_{\tau}$.
@@
For example, perhaps the simplest and most common parametric model for changepoint detection is the normal (Gaussian) meanshift model, in which the time series $y$ is assumed to be generated by a different random variable $Y_{j}\sim\mathcal{N}(\mu_{j},\sigma^{2})$ for each region $1\leq j\leq m+1$. Thus the series mean $\mu_{j}$ is constant within each of the $m+1$ regions, but varies across the regions (piecewise stationarity), while the variance $\sigma^{2}$ is constant over the entire time series. We fit the normal meanshift model using Maximium Likelihood Estimation (MLE), resulting in the estimated parameters $\hat{\theta}=\{\hat{\sigma}^{2},\hat{\mu}_{0},\ldots,\hat{\mu}_{m}\}$. If we choose the Bayesian Information Criterion (BIC) (Schwarz 1978) as our penalty, then the penalized objective function becomes:
@@
where $k_{M}(\tau)$ is the number of parameters estimated by the model (in this case, $m+2$) and $L_{M}(y|\hat{\theta}_{\tau})$ is the likelihood function for $M$, which in this case is (see S. Li and Lund (2012) for details):
@@
More generally, given a time series $y$, a parametric model $M(y|\theta)$, and a penalized objective function $f$, the goal of any changepoint detection algorithm $A$ is to search the exponentially large space of possible changepoints sets for the one $\tau^{*}$ that minimizes the value of $f(\tau,M(y|\hat{\theta}_{\tau}))$. Of course, in the process it must also estimate $\hat{\theta}_{\tau}$.
1.5 Selected changepoint detection algorithms
@@
The Pruned Exact Linear Time (PELT) algorithm developed by Killick, Fearnhead, and Eckley (2012) builds on a dynamic programming algorithm from Jackson et al. (2005) to offer a linear time, exact algorithm for changepoint detection. PELT offers order of magnitude improvements in computational cost (i.e., $O(n)$ vs. the $O(n\log{n})$ performance of the binary segmentation algorithm (Scott and Knott 1974) and the $O(n^{3})$ performance of the segmented neighborhood algorithm (Auger and Lawrence 1989)), and is capable of optimizing over different parametric models (e.g., meanvar, meanshift, etc.) and different penalized objective functions (e.g., BIC, MBIC (see Section 6.1.2), etc.). To guarantee an optimal solution, PELT requires that the value of the penalized objective function decreases (or increases by an amount bounded by a known constant) with each additional changepoint. Moreover, its running time can be quadratic (i.e., $O(n^{2})$) unless the number of changepoints grows linearly with the length of the time series. While these assumptions are mild, not all possible combinations of time series, models, and penalty functions will meet them. [^3] To see why, note that PELT’s reliance on dynamic programming requires that the penalized objective function be additive across the regions. Some penalty functions, such as MDL (see Section 6.1.3) depend on the locations of the changepoints, and thus cannot be additive across the regions.
@@
S. Li and Lund (2012) illustrate how a genetic algorithm can provide good results by employing Darwinian principles (e.g., survival of the fittest) to search the space of all possible changepoint sets intelligently. For a given parametric model and penalized objective function, a genetic algorithm starts with an initial “generation” of (usually randomly generated) candidate changepoint sets, evaluates them all using the penalized objective function, then probabilistically “mates” the most promising candidate changepoint sets to produce a new generation of the same size. This process is repeated until a stopping criteria is met, and the single best (fittest) changepoint set is returned. While genetic algorithms are neither exact, nor deterministic, nor computationally efficient, they are more easily generalizable than PELT, and can return competitive results in practice (Shi et al. 2022). S. Li and Lund (2012) also develop the notion of the Minimum Descriptive Length (MDL) as a penalty function based on information theory.
Bai and Perron (1998) use linear regression models fit by least squares to detect changepoints. In particular, they construct a framework for testing the alternative hypothesis that there is one additional changepoint. Zeileis et al. (2003) build on this and related work by Bai and Perron (2003), using a series of F-statistics and a dynamic programming algorithm to find the set of changepoints that minimize the residual sum of squared errors (see Section 1.6). V. M. Muggeo (2003) also employs a linear regression framework, but iteratively fit models using the Davies test (Davies 1987) as a criterion.
@@
Barry and Hartigan (1993) employ Bayesian techniques for changepoint detection. Cho and Fryzlewicz (2015) propose the Sparsified Binary Segmentation (SBS) algorithm, which uses CUSUM statistics to analyze high dimensional time series. This algorithm uses the concept of a threshold above which statistics are kept—an idea that resurfaces in Suárez-Sierra, Coen, and Taimal (2023). Cho (2016) explores the utility of a double CUSUM statistic for panel data. Hocking et al. (2013) use machine learning techniques to determine the optimal number of changepoints.
@@ 1.6 Existing changepoint detection R packages
There are dozens of changepoint packages on CRAN, all of which have their strengths and limitations. Please see the “Comparison to other packages” vignette in the mcp (Lindeløv 2024) documentation for a thorough comparison.
According to statistics returned by the pkgsearch (Csárdi and Salmon 2025) package, only three changepoint detection packages have more than 1000 downloads per month and more than two reverse dependencies: strucchange (Zeileis et al. 2024), segmented (V. M. R. Muggeo 2025), and changepoint (Killick 2024; Killick and Eckley 2014). tidychangepoint now supports all three.
strucchange (Zeileis et al. 2024, 2002) uses a regression-based framework and implements a dynamic programming algorithm to identify changepoints. The segmented (V. M. R. Muggeo 2025) package iteratively searches for optimal changepoint sets in a linear regression framework.
changepoint (Killick 2024; Killick and Eckley 2014) implements the algorithms PELT, Binary Segmentation, and Segmented Neighborhood. PELT employs one of three different models: meanshift, meanvar, and varshift (see Section 5). For the meanvar model (the most versatile) the data-generating model can follow a Normal, Poisson, Gamma, or Exponential distribution. Penalty functions (see Section 6) implemented by PELT include AIC, BIC, and MBIC, but not MDL. EnvCpt (Killick et al. 2025) is another package that wraps changepoint functionality.
The wbs (Baranowski and Fryzlewicz 2024) package implements the Wild Binary Segmentation and standard Binary Segmentation algorithms, the first of which is incorporated into tidychangepoint. The ggchangepoint (Yu 2022) package provides some wrapper functions and graphical tools for changepoint detection analysis that are similar to ours, but more limited in scope and usability. Nonparametric methods for changepoint detection are present in the ecp (James, Zhang, and Matteson 2024) and changepoint.np (Haynes and Killick 2022) packages. We consider these beyond our scope at this time, which is focused on parameteric models for changepoint detection.
A Bayesian approach is implemented in the bcp (X. Wang, Erdman, and Emerson 2018; Erdman and Emerson 2008) package, which is no longer available on CRAN. The qcc (Scrucca 2017) package is more broadly focused on quality control and statistical process control, neither of which are specific to changepoint detection. The mcp (Lindeløv 2024) package focuses on regression-based, Bayesian methods.
The changepointGA (M. Li and Lu 2025) package appeared on CRAN subsequent to the submission of this paper. It provides a detailed implementation of several changepoint detection algorithms that are built on top of a custom algorithmic implementation (M. Li and Lu 2024). Note that in contrast, tidychangepoint relies on a more general-purpose genetic algorithm from the GA (Scrucca 2024) package as a backend. We have included basic support for the ARIMA model with a BIC penalty function in tidychangepoint.